# Text-Free Multi-domain Graph Pre-training: Toward Graph Foundation Models

### v3

@@ C Further Descriptions of Baselines
In this section, we present more details for the baselines used in our experiments.
@@
  • GCN [8]: GCN utilizes a mean-pooling strategy for neighborhood aggregation to integrate information from neighboring nodes.
  • GAT [9]: GAT also leverages neighborhood aggregation for end-to-end node representation learning, uniquely assigns varying attention weights to different neighbors, thereby adjusting their impact on the aggregation process.
Table 4: Summary of datasets.
@@
  • DGI [9]: DGI operates as a self-supervised pre-training methodology tailored for homogeneous graphs. It is predicated on the maximization of mutual information (MI), aiming to enhance the estimated MI between locally augmented instances and their global counterparts.
  • InfoGraph [50]: Expanding upon DGI, InfoGraph is centered on graph-level tasks, endeavoring to maximize the alignment between node and graph embeddings.
  • GraphCL [20]: GraphCL leverages a variety of graph augmentations for self-supervised learning, tapping into the intrinsic structural patterns of graphs. The overarching goal is to amplify the concordance between different augmentations throughout graph pre-training.
@@
  • Hassani [25]: Hassani proposes an attention-based graph encoder that utilizes contextual and topological views of graph to learn task-specific knowledge for rapid adaptation and task-agnostic knowledge for knowledge transfer.
@@
  • GPPT [40]: GPPT utilize a GNN model, pre-trained through a link prediction task. The downstream prompt module is specifically tailored for node classification task, which is unified as the pre-training link prediction task.
  • GPF [51]: GPF operates as a universal prompt-based tuning methodology tailored for pre-trained GNN models. It is predicated on the adaptation of the input graph’s feature space, aiming to achieve an effect equivalent to any form of prompting function.
  • GraphPrompt [21]: GraphPrompt employs subgraph similarity calculation as a unified template to narrow the gap between pre-training and downstream tasks, inclusive of node and graph classification. A learnable prompt is subsequently tuned during the downstream adaptation to incorporate task-specific knowledge.
@@
  • GCOPE [33]: GCOPE is a multi-domain pre-training approach that cooperates graph datasets across multiple domains through a series of domain-specific interconnect virtual nodes that link nodes within the same domain. Its primary goal is to improve performance in downstream applications by leveraging the multiple domain knowledge in multiple source domains.
D Implementation Details
@@
Table 5: Settings of multi-domain transfer to unseen domains.
@@
For the baseline GCN [8], we employ a 3-layer architecture, and set the hidden dimensions to 256. For GAT [9], we employ a 2-layer architecture and set the hidden dimension to 64. Additionally, we apply 8 attention heads in the first GAT layer.
For DGI [9], we utilize a 1-layer GCN as the base model and set the hidden dimensions to 256. Additionally, we employ prelu as the activation function. For InfoGraph [50], a 3-layer GCN is used as the base model, with its hidden dimensions set to 256. For GraphCL [20], a 1-layer GCN is also employed as its base model, with the hidden dimensions set to 256. Specifically, we select edge dropping as the augmentations, with a default augmentation ratio of 0.2.
For Hassani [25],We employ a 3-layer GCN is used as the base model for all datasets, with the hidden dimensions set to 256.
For GPPT [40], we utilize a 2-layer GraphSAGE as its base model, setting the hidden dimensions to 256. For base GraphSAGE, we also employ a mean aggregator. For GraphPrompt [21], a 3-layer GCN is used as the base model for all datasets, with the hidden dimensions set to 256. For GPF [51], employs a 5-layer GCN as the base model for all datasets, following the recommended settings. The hidden dimensions are set to 256.
For GCOPE [33], we employ a 2-layer GCN as the base model and set the hidden dimensions to 100. Downstream adaptation is achieved through fine-tuning, as it is reported to yield the best performance in their literature.
@@
E Further Experiments and Analysis
Node classification on seen domains. We further conduct one-shot NC tasks on seen domains. Specifically, we pre-train jointly on Cora, Citeseer, Pubmed, Photo and Computers, and evaluate the NC tasks on Cora and Citeseer. We illustrate the results of competitive baselines and MDGPT in Table 6, and observe the same trend as those observed in in Sect. 5.2, showing the robustness of MDGPT in multi-domain transfer setting.
Table 6: Accuracy of one-shot node classification on seen target domains.
MethodTarget domain Cora Citeseer
DGI 33.57 $\pm$ 06.83 32.90 $\pm$ 06.10
GraphCL 38.73 $\pm$ 07.43 32.46 $\pm$ 04.75
GraphPrompt 45.98 $\pm$ 09.43 36.45 $\pm$ 08.38
GCOPE 34.95 $\pm$ 07.45 35.18 $\pm$ 07.76
MDGPT 47.44 $\pm$ 10.03 44.96 $\pm$ 10.12
Impact of shots on graph classification. We vary the number of shots and conduct graph classification tasks on three target domains, i.e., Cora, Photo and Reddit. The results are illustrated in Fig. 6, and we observe same patterns as shown in node classification tasks.
Figure 6: Impact of shots on graph classification.
Table 7: Comparison of the number of tunable parameters during the downstream adaptation stage.
MethodTarget domain Cora Citeseer Pubmed Photo Computers Reddit
GCN 14,592 14,336 13,568 14,848 15,360 23,296
GraphCL 1,792 1,536 768 2,048 2,560 10,496
GraphPrompt 256 256 256 256 256 256
GCOPE 1,792 1,536 768 2,048 2,560 10,496
MDGPT 56 56 56 56 56 57
Parameters efficiency. We evaluate the parameter efficiency of MDGPT compared to representative baselines. In particular, we calculate the number of tunable parameters during the downstream adaptation stage, and present the results in Table 7. For GCN, as it follows an end-to-end supervised paradigm, all parameters in the model have to be updated, resulting in the worst parameter efficiency. For GraphCL and GCOPE, we update only the downstream classifier while frozen the pre-trained model, signficantly reduce the number of tunable parameters. For GraphPrompt, since it only tunes a light-wight prompt containing fewer parameters than a classifier (e.x., fully connected layer), the number of parameters is further reduced. Finally, MDGPT is the most parameter-efficient adaptation, as only the dual prompts are updated. Specifically, the unifying prompt has the same dimension as unified feature, which is generally smaller than embedding dimensions, and the mixing prompt’s dimension is same as the number of source domains, which is a quite small compared to feature dimension. Despite having the fewest parameters, MDGPT significantly outperforms all baselines, further demonstrating the effectiveness of our design.

### v4

@@ B. Further Descriptions of Baselines
In this section, we summarize the datasets in Table 7 and present more details for the baselines used in our experiments.
@@
  • GCN (Kipf and Welling, 2017): GCN utilizes a mean-pooling strategy for neighborhood aggregation to integrate information from neighboring nodes.
  • GAT (Veličković et al., 2018a): GAT also leverages neighborhood aggregation for end-to-end node representation learning, uniquely assigns varying attention weights to different neighbors, thereby adjusting their impact on the aggregation process.
Table 7. Summary of datasets.
@@
  • DGI (Veličković et al., 2018a): DGI operates as a self-supervised pre-training methodology tailored for homogeneous graphs. It is predicated on the maximization of mutual information (MI), aiming to enhance the estimated MI between locally augmented instances and their global counterparts.
  • InfoGraph (Sun et al., 2019): Expanding upon DGI, InfoGraph is centered on graph-level tasks, endeavoring to maximize the alignment between node and graph embeddings.
  • GraphCL (You et al., 2020): GraphCL leverages a variety of graph augmentations for self-supervised learning, tapping into the intrinsic structural patterns of graphs. The overarching goal is to amplify the concordance between different augmentations throughout graph pre-training.
@@
  • Hassani (Hassani, 2022): Hassani proposes an attention-based graph encoder that utilizes contextual and topological views of graph to learn task-specific knowledge for rapid adaptation and task-agnostic knowledge for knowledge transfer.
@@
  • GPPT (Sun et al., 2022): GPPT utilize a GNN model, pre-trained through a link prediction task. The downstream prompt module is specifically tailored for node classification task, which is unified as the pre-training link prediction task.
  • GPF (Fang et al., 2024): GPF operates as a universal prompt-based tuning methodology tailored for pre-trained GNN models. It is predicated on the adaptation of the input graph’s feature space, aiming to achieve an effect equivalent to any form of prompting function.
  • GraphPrompt (Liu et al., 2023b): GraphPrompt employs subgraph similarity calculation as a unified template to narrow the gap between pre-training and downstream tasks, inclusive of node and graph classification. A learnable prompt is subsequently tuned during the downstream adaptation to incorporate task-specific knowledge.
@@
  • GCOPE (Zhao et al., 2024): GCOPE is a multi-domain pre-training approach that cooperates graph datasets across multiple domains through a series of domain-specific interconnect virtual nodes that link nodes within the same domain. Its primary goal is to improve performance in downstream applications by leveraging the multiple domain knowledge in multiple source domains.
C. Implementation Details
@@
Table 8. Settings of multi-domain transfer to unseen domains.
@@
For the baseline GCN (Kipf and Welling, 2017), we employ a 3-layer architecture, and set the hidden dimensions to 256. For GAT (Veličković et al., 2018a), we employ a 2-layer architecture and set the hidden dimension to 64. Additionally, we apply 8 attention heads in the first GAT layer.
For DGI (Veličković et al., 2018a), we utilize a 1-layer GCN as the base model and set the hidden dimensions to 256. Additionally, we employ prelu as the activation function. For InfoGraph (Sun et al., 2019), a 3-layer GCN is used as the base model, with its hidden dimensions set to 256. For GraphCL (You et al., 2020), a 1-layer GCN is also employed as its base model, with the hidden dimensions set to 256. Specifically, we select edge dropping as the augmentations, with a default augmentation ratio of 0.2.
For Hassani (Hassani, 2022),We employ a 3-layer GCN is used as the base model for all datasets, with the hidden dimensions set to 256.
For GPPT (Sun et al., 2022), we utilize a 2-layer GraphSAGE as its base model, setting the hidden dimensions to 256. For base GraphSAGE, we also employ a mean aggregator. For GraphPrompt (Liu et al., 2023b), a 3-layer GCN is used as the base model for all datasets, with the hidden dimensions set to 256. For GPF (Fang et al., 2024), employs a 5-layer GCN as the base model for all datasets, following the recommended settings. The hidden dimensions are set to 256.
For GCOPE (Zhao et al., 2024), we employ a 2-layer GCN as the base model and set the hidden dimensions to 100. Downstream adaptation is achieved through fine-tuning, as it is reported to yield the best performance in their literature.